# A Simple LLM Framework for Long-Range Video Question-Answering

### v1

@@ 3.3 Implementation Details
For the experiments on EgoSchema, we use LaViLa [83] as our short-term captioner. We note that LaViLa is pretrained on Ego4D, the same dataset that the authors of EgoSchema use to build their benchmark. Thus, to avoid any data intersection between training and testing sets, we retrain our variant of the LaViLa model on a subset of 6K Ego4D videos that do not include any EgoSchema videos. To process video inputs, we segment each video into multiple 1s clips with a stride of 1s, resulting in a list of consecutive clips that cover the entire video. We use GPT-3.5 as the LLM for long-range reasoning on EgoSchema. For NeXT-QA, IntentQA and NeXT-GQA, we use LLaVA-1.5 [36] as the visual captioner and GPT-4 as the LLM. We downsample the videos to 0.5 FPS and prompt LLaVA to generate captions that contain roughly 30 words for each frame. We use LaViLa as our visual captioner on EgoSchema because compared to other captioners (e.g. BLIP-2, LLaVA), LaViLa works better on first-person view videos. For the third-person view video datasets, we use LLaVA because we found its performance to be better than LaViLa. We also observe that GPT-4 generally performs better than other language models. However, on EgoSchema, we found the output of GPT-4 inconclusive for many questions (e.g. an output message that the provided information is insufficient to answer the given question). In contrast, we did not experience such issues with GPT-3.5. We also did not observe these issues with GPT-4 on other datasets except EgoSchema. We provide more implementation details in the Supplementary Material.
@@ 4 Experiments
Figure 4: An illustration of prior LVQA dataset/benchmark limitations. Top: An example from MovieQA [52]. The model can use the provided subtitle information to answer a question while completely ignoring visual cues in a video. Middle: An example from ActivityNet-QA Dataset [80]. Despite long video inputs, the model only needs to analyze a short 1s video clip to answer the question. Bottom: An example from EgoShcema Dataset [37]. The model must analyze visual cues from the whole 3-minute video to answer a given question without relying on any additional textual inputs (e.g., speech, subtitles).
@@ 4.1 Datasets and Metrics
Unlike short-term video question-answering, long-range video question-answering (LVQA) lacks robust and universally agreed-upon benchmarks. Many prior movie-based long-range video understanding benchmarks [52, 8] exhibit significant language biases, i.e., pure text-based approaches can achieve excellent performance by using information from subtitles/speech while ignoring the video content entirely [75, 21] (See the example in the top row of Figure 4). Furthermore, while many existing long-range video understanding benchmarks [80, 65, 38] contain long video inputs, prior work [37] have shown that all of these benchmarks can be solved by analyzing only several second-long video clips within the longer video inputs, without requiring any long-range video modeling capabilities (See the example in the middle row of Figure 4).
To address these limitations, recent work introduced EgoSchema [37], a new long-range video question-answering benchmark, consisting of 5K multiple choice question-answer pairs, spanning 250 hours of video, and covering a wide range of human activities. Unlike prior benchmarks in this area, the authors in [37] have manually verified that to answer a given question correctly, one must watch at least 100 seconds of video content, which is orders of magnitude longer than any existing benchmark (See the example in the bottom row of Figure 4). Furthermore, unlike prior LVQA benchmarks, EgoSchema has no textual inputs such as speech, subtitles, or storylines. This means the questions must be answered based only on video inputs, preventing language-based biases. The EgoSchema benchmark consists of 5,000 questions, each requiring the correct answer to be selected between 5 given options based on a three-minute-long video clip. The entire dataset is designed for zero-shot evaluation and has no training set. For validation, the authors released a subset of 500 questions with ground truth answers (EgoSchema Subset). By default, our experiments are conducted on the EgoSchema Subset. The metric we use is QA accuracy, i.e., the percentage of correctly answered questions among all questions.
In addition to EgoSchema, we also perform zero-shot LVQA experiments on three other LVQA benchmarks:
  • NExT-QA [68] contains 5,440 videos with an average duration of 44s and 48K multi-choice questions and 52K open-ended questions. There are 3 different question types: Temporal, Causal, and Descriptive. Following common practice, we perform zero-shot evaluation on the validation set, which contains 570 videos and 5K multiple-choice questions.
  • IntentQA [30] contains 4,303 videos and 16K multiple-choice question-answer pairs focused on reasoning about people’s intent in the video. We perform a zero-shot evaluation on the test set containing 2K questions.
  • NExT-GQA [71] is an extension of NExT-QA with 10.5K temporal grounding annotations associated with the original QA pairs. The dataset was introduced to study whether the existing LVQA models can temporally localize video segments needed to answer a given question. We evaluate all methods on the test split, which contains 990 videos with 5,553 questions, each accompanied by a temporal grounding label. The metrics we used include: 1) Intersection over Prediction (IoP) [71], which measures whether the predicted temporal window lies inside the ground truth temporal segment, 2) temporal Intersection over Union (IoU), and 3) Acc@GQA, which depicts the percentage of accurately answered and grounded predictions. For IoP and IoU, we report the mean values and values with the overlap thresholds of 0.5.
To study the most important factors in our framework design, we first conduct an empirical study on the released subset of EgoSchema. Afterward, we present our main results on the full EgoSchema test set. Lastly, we present our results on other datasets and tasks.
@@ 4.2 Empirical Study on EgoSchema
Before presenting our main results, we first study the effectiveness of different components within our LLoVi framework, including (i) the visual captioner, (ii) the LLM, (iii) the LLM prompt design, and (iv) few-shot in-context learning. The experiments are conducted on the EgoSchema Subset with 500 multi-choice questions. We discuss our empirical findings below. We also include additional experiments in the supplementary material.
@@
VideoBLIP clip-level 40.0
EgoVLP clip-level 46.6
BLIP-2 frame-level 46.7
LaViLa clip-level 51.8
Oracle clip-level - 65.8
@@
4.2.1 Visual Captioning Model
In Table 1, we study the effectiveness of various clip-level video captioners, including LaViLa [83], EgoVLP [43], and VideoBLIP [78]. In addition to video captioners, we also try the state-of-the-art image captioner, BLIP-2 [28]. Lastly, to study the upper bound of our visual captioning results, we include the ground truth Oracle captioning baseline obtained from the Ego4D dataset. All baselines in Table 1 use similar experimental settings, including the same LLM model, i.e., GPT-3.5. The results are reported as LVQA accuracy on the EgoSchema Subset.
Based on the results in Table 1, we observe that LaViLa is the best captioning model, outperforming BLIP-2, EgoVLP, and VideoBLIP. We also observe that despite not being pre-trained on Ego4D [12], BLIP-2 performs reasonably well (46.7%) and even outperforms other strong Ego4D-pretrained baselines, EgoVLP and VideoBLIP. Lastly, the Oracle baseline with ground truth captions outperforms LaViLa captions by a large margin (14.0%). This demonstrates that our framework can benefit from future improvements in visual captioning models.
@@ 4.2.1 Visual Captioning Model
Llama2-7B 7B 34.0
Llama2-13B 13B 40.4
Llama2-70B 70B 50.6
GPT-3.5 175B 51.8
GPT-4 N/A 58.3
Table 2: Accuracy of our framework with different LLMs. GPT-4 achieves the best accuracy, suggesting that stronger LLMs perform better in LVQA. However, due to the best accuracy and cost tradeoff, we use GPT3.5 for our remaining experiments.
@@ 4.2.2 Large Language Model
In Table 2, we analyze the performance of our LLoVi framework using different LLMs while fixing the visual captioner to be LaViLa. Based on these results, we observe that GPT-4 achieves the best performance (58.3%), followed by GPT-3.5 (51.8%). These results suggest that stronger LLMs (GPT-4) are better at long-range modeling, as indicated by a significant margin in LVQA accuracy between GPT-4 and all other LLMs ($>$6.5%). We also note that the Llama2 performs reasonably well with its 70B variant (50.6%), but its performance drastically degrades with smaller capacity LLMs (i.e., Llama2-7B, Llama2-13B). Due to the tradeoff between accuracy and cost, we use GPT-3.5 for all of our remaining experiments unless noted otherwise.
@@ 4.2.3 LLM Prompt Analysis
In this section, we (1) analyze several variants of our summarization-based prompt (described in Section 3), and (2) experiment with other commonly used prompt designs, including Zero-shot Chain-of-Thought (Zero-shot CoT) [64], Plan-and-Solve [57], and Self-Consistency [59]. Below, we present a detailed analysis of these results.
A Multi-round Summarization-based Prompt. As discussed in Section 3, we found that a specialized multi-round prompt that first asks the LLM to summarize the noisy short-term captions, and then answer the question using the LLM-generated summary performs better than our base approach that uses caption inputs directly for LVQA. Given a concatenated set of captions $C$, an input question $Q$, and a set of candidate answers $A$, there are several input combinations that we can use to obtain the summary $S$. Thus, here, we investigate three distinct variants of obtaining summaries $S$:
  1. 1.
@@
  3. 3.
    (C, Q, A) $\rightarrow$ S: the LLM takes captions $C$, a question $Q$, and the answer candidates $A$ as its inputs to produce summaries $S$. Like above, having additional answer candidate inputs enables the LLM to generate a summary $S$ most tailored to particular question-answer pairs.
Variant Standard (C) $\rightarrow$ S (C, Q) $\rightarrow$ S (C, Q, A) $\rightarrow$ S
Acc. (%) 51.8 53.6 57.6 55.9
Table 3: Different variants of our multi-round summarization-based prompt. Our results indicate that the (C, Q) normal-$\rightarrow$ S variant that takes concatenated captions $C$ and a question $Q$ for generating a summary $S$ works the best, significantly outperforming (+5.8%) the standard prompt. This confirms our hypothesis that additional inputs in the form of a question $Q$ enable the LLM to generate a summary $S$ tailored to a given question $Q$.
Number of words 50 100 300 500 700
Acc. (%) 55.6 57.4 55.8 57.6 55.0
Table 4: Number of words in a generated summary. We study the optimal number of words in an LLM-generated summary. These results suggest that the optimal LVQA performance is obtained when using 500-word summaries.
In Table 3, we explore the effectiveness of these three prompt variants. Our results show that all three variants significantly outperform our simple, yet already strong baseline that uses a standard LVQA prompt (described in Section 3). Specifically, we note that the variant (C) $\rightarrow$ S that uses caption-only inputs to obtain the summaries outperforms the standard baseline by 1.8%. Furthermore, we observe that incorporating a given question as an input (i.e., the (C, Q) $\rightarrow$ S variant) leads to the best performance (57.6%) with a significant 5.8% boost over the standard LVQA prompt baseline. This confirms our earlier intuition that having additional question $Q$ inputs enables the LLM to generate a summary $S$ specifically tailored for answering that question, thus leading to a big boost in LVQA performance. Lastly, we observe that adding answer candidates $A$ as additional inputs (i.e., the (C, Q, A) $\rightarrow$ S variant) leads to a drop in performance (-1.7%) compared with the (C, Q) $\rightarrow$ S variant. We conjecture that this might be because the wrong answers in the candidate set $A$ may mislead the LLM, leading to a suboptimal summary $S$.
We also investigate the optimal length of the generated summary $S$, and present these results in Table 4. Specifically, for these experiments, we ask the LLM to generate a summary $S$ using a different number of words (as part of our prompt). We use the best performing (C, Q) $\rightarrow$ S variant for these experiments. Our results indicate that using a very small number of words (e.g., 50) leads to a drop in performance, indicating that compressing the caption information too much hurts the subsequent LVQA performance. Similarly, generating summaries that are quite long (e.g., 700 words) also leads to worse results, suggesting that the filtering of the potentially noisy/redundant information in the captions is important for good LVQA performance. The best performance is obtained using 500-word summaries.

### v3

@@ 3.3 Implementation Details
For the experiments on EgoSchema, we use LaViLa Zhao et al. (2023) as our captioner. We segment each video into multiple 1s clips, resulting in a list of consecutive clips that cover the entire video. We use GPT-3.5 as the LLM on EgoSchema. For NExT-QA, IntentQA, and NExT-GQA, we use CogAgent Hong et al. (2024) as the visual captioner and GPT-4 as the LLM. We downsample the videos to 0.5 FPS and prompt CogAgent to generate captions for each frame. More details are provided in the Supplementary Material.
@@ 4 Experiments
Figure 4: An illustration of prior LVQA dataset limitations. Top: An example from MovieQA Tapaswi et al. (2016). The model can use the provided subtitle information to answer a question while ignoring visual cues in a video. Middle: An example from the ActivityNet-QA Dataset Yu et al. (2019). Despite long video inputs, the model only needs to analyze a short 1s video clip to answer the question. Bottom: An example from the EgoSchema Dataset Mangalam et al. (2023). The model must analyze visual cues from the video to answer a given question without relying on additional textual inputs (e.g., speech, subtitles).
@@ 4.1 Datasets and Metrics
Unlike short-term video question-answering, long-range video question-answering (LVQA) lacks robust and universally agreed-upon benchmarks. As shown in Figure 4, many prior LVQA benchmarks either exhibit significant language biases, or do not require long-range video modeling capabilities. To address these limitations, recent work introduced EgoSchema Mangalam et al. (2023), a new long-range video question-answering benchmark consisting of 5K multiple choice question-answer pairs spanning 250 hours of video and covering a wide range of human activities. By default, our experiments are conducted on the validation set of 500 questions (referred to as the EgoSchema Subset). The final comparison is done on the full test set of 5K EgoSchema questions. We use QA accuracy (i.e., the percentage of correctly answered questions) as our evaluation metric. Additionally, we also perform zero-shot LVQA experiments on three commonly-used LVQA benchmarks: NExT-QA Xiao et al. (2021), IntentQA Li et al. (2023a), and NExT-GQA Xiao et al. (2023). Detailed dataset information and metrics can be found in the Supplementary Material.
@@ 4.2 Empirical Study on EgoSchema
We first study the effectiveness of different components within our LLoVi framework, including (i) the visual captioner, (ii) the LLM, (iii) the optimal video processing configurations, and (iv) the LLM prompt design. The experiments are conducted on the EgoSchema Subset. We discuss our empirical findings below. We also include additional experiments in the Supplementary Material.
4.2.1 Visual Captioning Model
@@
EgoVLP Qinghong Lin et al. (2022) clip-level 46.6
LLaVA Liu et al. (2023b) frame-level 45.2
BLIP-2 Li et al. (2023d) frame-level 50.6
LaViLa Zhao et al. (2023) clip-level 55.2
Oracle clip-level - 66.0
@@
In Table 1, we study the effectiveness of various clip-level video captioners, including LaViLa Zhao et al. (2023) and EgoVLP Qinghong Lin et al. (2022). In addition to video captioners, we also try the state-of-the-art image captioners, BLIP-2 Li et al. (2023c) and LLaVA-1.5 Liu et al. (2023b). Lastly, to study the upper bound of our visual captioning results, we include the ground truth Oracle captioning baseline obtained from the Ego4D dataset. All baselines in Table 1 use similar experimental settings, including the same LLM model, i.e., GPT-3.5. The results are reported as LVQA accuracy on the EgoSchema Subset. Table 1 suggests that LaViLa provides the best results, outperforming BLIP-2, EgoVLP, and LLaVA. We also observe that despite not being pre-trained on Ego4D Grauman et al. (2022), BLIP-2 performs reasonably well (50.6%) and even outperforms a strong Ego4D-pretrained baseline, EgoVLP. Lastly, the Oracle baseline with ground truth captions outperforms LaViLa captions by a large margin (10.8%). This shows that our method can benefit from future improvements in captioning models.
In addition to our quantitative analysis, we also observed that our framework with the LaViLa captioner demonstrates basic Person Re-Identification capabilities when the video involves simple interactions among people. We visualize these results in our Supplementary Material.
@@
Mistral Jiang et al. (2023) 7B 50.8
Llama3-8B Touvron et al. (2023) 8B 52.2
Llama3-70B Touvron et al. (2023) 70B 56.8
GPT-3.5 OpenAI (2023a) 175B 55.2
GPT-4 OpenAI (2023b) 1.8T 61.2
Table 2: Accuracy of our framework with different LLMs. GPT-4 achieves the best accuracy, suggesting that stronger LLMs perform better in LVQA. However, we use GPT-3.5 for most of our experiments due to the best accuracy and cost tradeoff.
@@ 4.2.2 Large Language Model
In Table 2, we analyze the performance of our framework using different LLMs while fixing the visual captioner to be LaViLa. Our results indicate that GPT-4 achieves the best performance (61.2%), followed by LLama3-70B (56.8%) and GPT-3.5 (55.2%). Thus, stronger LLMs (GPT-4) are better at long-range modeling, as indicated by a significant margin in LVQA accuracy between GPT-4 and all other LLMs ($>$4.4%). We also observe that despite having a much smaller number of parameters, LLama3-8B (52.2%) and Mistral-7B (50.8%) still achieve competitive performance. Due to the high cost of GPT-4 and the large computational resource requirements of Llama3-70B, we use GPT-3.5 for most of our experiments unless noted otherwise.
Clip length (s) 1 2 4 8
Acc. (%) 55.2 54.8 53.4 53.4
Table 3: Analysis of different clip length. We divide the input long video into consecutive clips of different length. The highest accuracy is achieved when the clips are shortest, while performance diminishes as clip length increases. This indicates that splitting long videos into shorter segments, particularly 1-second clips, is the most efficient approach.
Clip sampling rate 1 1/2 1/4 1/8
Acc. (%) 55.2 55.2 54.6 53.2
Table 4: Analysis of sparse video clip sampling. We divide the input long video into consecutive 1s short clips and study the effect of different clip sampling rates. Sampling clips every 1s achieves the best performance while sampling clips every 8s achieves the best efficiency (8x) with only 2.0% accuracy drop. This suggests that we can effectively control the accuracy-efficiency trade-off of our framework by varying the clip sampling rate.
4.2.3 Video Processing Configurations
Clip length and sample rate are important hyper-parameters for sampling short video clips from long video inputs for visual captioning. In this section, we explore the influence of clip length and clip sampling rate on our framework.
Clip Length. In Table 3, we explore how LVQA performance is influenced by different clip length. We divide the long video into consecutive clips of different length and report the corresponding LVQA accuracy. From the table, we can see that our framework achieves the best accuracy when the clip length is the shortest. As the clip length increases, the performance drops. This suggests that dividing long videos into consecutive 1s short clips is the most effective strategy.
Clip Sampling Rate. In Table 4, we explore how LVQA performance is influenced by different clip sampling rate on EgoSchema. Specifically, we divide the input long video into consecutive 1s short clips and change the clip sampling rate to see how LVQA performance changes accordingly. From the table, we can see that sampling one clip every 1s leads to the highest accuracy. Sampling one clip every 8s (i.e., the clip sampling rate of 1/8) achieves 8x efficiency while the accuracy drops by only 2.0%. This indicates that we can effectively control the accuracy and efficiency tradeoff of our method by sampling video clips more sparsely.
@@ 4.2.4 LLM Prompt Analysis
In this section, we (1) analyze several variants of our summarization-based prompt (described in Section 3), and (2) experiment with other commonly used prompt designs, including Zero-shot Chain-of-Thought (Zero-shot CoT) Wei et al. (2022) and Plan-and-Solve Wang et al. (2023c). Below, we present a detailed analysis of these results.
Multi-round Summarization Prompt. Given a concatenated set of captions $C$, an input question $Q$, and a set of candidate answers $A$, we can use several input combinations to obtain the summary $S$. Thus, here, we investigate three distinct variants of obtaining summaries $S$:
@@
  • (C, Q, A) $\rightarrow$ S: the LLM takes captions $C$, a question $Q$, and the answer candidates $A$ as its inputs to produce summaries $S$. Having additional answer candidate inputs enables the LLM to generate a summary $S$ most tailored to particular question-answer pairs.
Prompt Type Standard (C) $\rightarrow$ S (C, Q) $\rightarrow$ S (C, Q, A) $\rightarrow$ S
Acc. (%) 55.2 55.0 58.8 54.8
Table 5: Different variants of our multi-round summarization prompt. Our results indicate that the (C, Q) $\rightarrow$ S variant that takes concatenated captions $C$ and a question $Q$ for generating a summary $S$ works the best, significantly outperforming (+3.6%) the standard prompt. This confirms our hypothesis that additional inputs in the form of a question $Q$ enable the LLM to generate a summary $S$ tailored to a given question $Q$.
In Table 5, we explore the effectiveness of these three prompt variants. We observe that while the (C) $\rightarrow$ S and the (C, Q, A) variant $\rightarrow$ S perform similarly to the standard baseline, the (C, Q) $\rightarrow$ S variant greatly outperforms the standard baseline by 3.6%. Compared with (C) $\rightarrow$ S, (C, Q) $\rightarrow$ S incorporates a given question as the input and thus leads to a big boost in LVQA performance. This confirms our earlier intuition that having additional question $Q$ inputs enables the LLM to generate a summary $S$ specifically tailored for answering that question. However, adding answer candidates $A$ as additional inputs (i.e., the (C, Q, A) $\rightarrow$ S variant) leads to a drop in performance (-4.0%) compared with the (C, Q) $\rightarrow$ S variant. We conjecture that this might happen because the candidate answers $A$ in EgoSchema are often very long, and thus, they may mislead/distract the LLM into generating a suboptimal summary $S$.
@@
      Standard       55.2
      Plan-and-Solve Wang et al. (2023c)       55.2
      Chain-of-Thought Wei et al. (2022)       57.8
      Ours       58.8
Table 6: Comparison with commonly used prompting techniques. The “Standard" means a standard LVQA prompt (see Section 3). Our multi-round summarization prompt performs best.
Comparison with Commonly Used Prompts. Next, in Table 6, we compare our multi-round summarization prompt with other commonly used prompts such as Zero-shot Chain-of-Thought Wei et al. (2022) and Plan-and-Solve Wang et al. (2023c). Our results indicate that our multi-round summarization prompt achieves the best performance among all of these prompts. Furthermore, we note that it outperforms the standard prompt (described in Section 3) by a substantial 3.6% in LVQA accuracy, thus indicating the effectiveness of our prompt design.
Efficiency Analysis. We compare the efficiency of our multi-round summarization prompt and the standard prompt within our entire framework. We report that for a 3-minute EgoSchema video, the LaViLa captioner takes 22.36s to generate all short-term captions on a single A6000 GPU. The standard prompt using GPT-3.5 as the LLM then takes 0.4s for processing the captions from the 3-minute video, while the multi-round summarization prompt takes 3.6s. Therefore, the additional computational cost introduced by the multi-round summarization prompt is relatively small compared to the total runtime, which shows the efficiency of our multi-round summarization prompt. We also note that such a small increase in runtime leads to a substantial 9.4% increase in QA accuracy on the full set of EgoSchema compared to using the standard prompt as shown in Table 7.
Method LM Params Throughput Acc. (%)
(video / s)
FrozenBiLM DeBERTa 900M - 26.9
mPLUG-Owl LLaMA 7B - 31.1
InternVideo Transformer 478M - 32.1
LongViViT BERT 1B - 33.3
Video ChatCaptioner GPT-3.5 175B 1.24 39.0
VLog GPT-3.5 175B 1.04 44.0
Vamos GPT-4 1.5T - 48.3
LLoVi (Ours) GPT-3.5 175B 2.63 42.8
w/ Standard Prompt
LLoVi (Ours) GPT-3.5 175B 2.31 52.2
w/ Summarization Prompt
Table 7: Main results on the full set of EgoSchema. The throughput is measured by the number of 3-minute videos that a method can process in one minute using an A6000 GPU. Our LLoVi framework with the proposed multi-round summarization prompt achieves 52.2% accuracy, outperforming the variant of our model with a standard prompt by a significant margin (9.4%). Additionally, our method outperforms the previous best-performing Vamos model by 3.9% despite using a weaker LLM, as well as all other competing methods. Our method also has the highest throughput compared with other LLM-based methods.
4.3 Main Results on EgoSchema
In Table 7, we evaluate our best-performing LLoVi framework on the full EgoSchema test set containing 5K video samples. We compare our approach with prior state-of-the-art methods including InternVideo Wang et al. (2022a), mPLUG-Owl Ye et al. (2023), FrozenBiLM Yang et al. (2022a), Video ChatCaptioner Chen et al. (2023), VLog Lin and Lei (2023), as well as the concurrent works of LongViViT Papalampidi et al. (2023), and Vamos Wang et al. (2023d). The throughput is measured by the number of 3-minute videos that a method can process in one minute using an A6000 GPU.
Based on these results, we observe that our LLoVi framework with the proposed multi-round summarization prompt achieves 52.2% accuracy, outperforming the concurrent Vamos model by +3.9% despite using a weaker LLM (GPT-3.5) than their approach (GPT-4). We also observe that our model outperforms all other baselines by an even more significant margin ($>$8.2%). Additionally, we can see that our method has the highest throughput compared with other LLM-based approaches. This shows that our framework is the most efficient while achieving the highest accuracy. Lastly, our results indicate that using our novel multi-round summarization prompt outperforms the variant of our approach with the standard prompt by a significant margin of 9.4%. These results validate the effectiveness of our LLM-based framework design.
4.4 Results on Other Datasets
Method LM Params C. T. D. All
VFC Transformer 164M 45.4 51.6 64.1 51.5
InternVideo Transformer 478M 43.4 48.0 65.1 49.1
ViperGPT GPT-3 175B - - - 60.0
SeViLA Flan-T5 4B 61.3 61.5 75.6 63.6
LLoVi (ours) GPT-3.5 175B 67.1 60.1 76.5 66.3
LLoVi (ours) GPT-4 1.8T 73.7 70.2 81.9 73.8
Table 8: Zero-shot results on NExT-QA. C, T, D is short for Causual, Temporal, Descriptive, respectively. The best variant of LLoVi achieves 73.8% accuracy, outperforming previous best-performing model SeViLA by 10.2%.
Method LM Params Acc. (%)
Supervised
HQGA - 46M 47.7
VGT Transformer 511M 51.3
BlindGPT GPT-3 175B 51.6
CaVIR GPT-3 175B 57.6